# Log-majorizations between quasi-geometric type means for matrices

### v2

@@ 5.1 Joint concavity/convexity and monotonicity
The next theorem was shown in [30, Theorem 3.2] in a more general form by a complex function method using Pick functions (called Epstein’s method).
@@
In particular, $\mathrm{Tr}\,\mathcal{A}_{\alpha,p}$, $\mathrm{Tr}\,\mathcal{H}_{\alpha,p}$ and $\mathrm{Tr}\,G_{\alpha,p}$ are jointly concave for $0<\alpha\leq 1$ and $0<p\leq 1$. More specified results on the joint concavity/convexity of $\mathrm{Tr}\,\mathcal{A}_{\alpha,p}$ and $\mathrm{Tr}\,\mathcal{H}_{\alpha,p}$ are known as follows (see [8, 15] and [30, p. 1583]):
@@
Let $0<\alpha<1$, $p>0$ and $n\in\mathbb{N}$, $n\geq 2$ be arbitrary. Then the following hold:
@@ 5.2 Quantum Rényi type divergences
where a POVM on $\mathbb{C}^{n}$ is a family $(M_{i})_{i=1}^{k}\subset\mathbb{M}_{n}^{+}$ with $\sum_{i=1}^{k}=I_{n}$ (i.e., a $k$-outcome positive operator-valued measure). Furthermore, the regularized measured $\alpha$-Rényi divergence is defined by
@@
we have the second inequality in (5.5). If, in addition, $D_{\alpha}^{\mathrm{q}}$ is additive under tensor products, then by definition (5.4) as well as (5.5) we furthermore have
@@
Let $p>0$. Either when $\mathcal{M}\in\{R,G,LG\}$ and $\alpha\in(0,\infty)\setminus\{1\}$, or when $\mathcal{M}\in\{SG,\widetilde{S}G\}$ and $\alpha>1$, the function $D^{\mathcal{M}_{\alpha,p}}$ defined in (5.7) is a quantum $\alpha$-Rényi type divergence in the sense of Definition 5.4, and is additive under tensor products.
@@
    ([45, Remark II.6]) When $\alpha>2$, $D_{\alpha}^{\max}(B\|A)\leq D^{G_{\alpha,1}}(B\|A)$ and $D_{\alpha}^{\max}\neq D^{G_{\alpha,1}}$.
@@
    ([46, Theorem 3.7], [27]) When $1/2\leq\alpha<\infty$,
@@
For any $\mathcal{M}_{\alpha,p}$ in Lemma 5.5, since $\mathcal{M}_{\alpha,p}(a,b)=a^{1-\alpha}b^{\alpha}$ in $a,b\in(0,\infty)$ is not convex when $0<\alpha<1$ and not concave when $\alpha>1$, we may consider only the joint concavity of $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ on $\mathbb{M}_{n}^{+}\times\mathbb{M}_{n}^{+}$ for $0<\alpha<1$, and the joint convexity of $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ on ($\mathbb{M}_{n}^{+}\times\mathbb{M}_{n}^{+})_{\geq}$ when $\alpha>1$. Below we will simply write $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ is jointly concave/convex” if it is so on its domain $\mathbb{M}_{n}^{+}\times\mathbb{M}_{n}^{+}$ or $(\mathbb{M}_{n}^{+}\times\mathbb{M}_{n}^{+})_{\geq}$ for any $n\in\mathbb{N}$, equivalently, if it is so on $\mathbb{M}_{n}^{++}\times\mathbb{M}_{n}^{++}$ for any $n\in\mathbb{N}$ (thanks to Proposition 2.1(2)).
@@
for all $(A,B)\in\mathrm{dom}\,\mathcal{M}_{\alpha,p}$, $B\neq 0$. When $\mathcal{M}\in\{SG,\widetilde{S}G\}$ and $0<\alpha<1$, even though $D^{\mathcal{M}_{\alpha,p}}$ is not precisely a quantum $\alpha$-Rényi type divergence in Definition 5.4, we can still have (5.8) for all $(A,B)\in\mathrm{dom}\,\mathcal{M}_{\alpha,p}$. In fact, Theorem 5.3 implies that $D^{\mathcal{M}_{\alpha,p}}$ is monotone under CPTP maps with restriction to $\mathrm{dom}\,\mathcal{M}_{\alpha,p}$. Hence by Lemma 5.6 the first inequality in (5.8) holds on $\mathrm{dom}\,\mathcal{M}_{\alpha,p}$. Also, note (see [44] and the proof of [35, Proposition 4.1]) that for any $(A,B)\in\mathrm{dom}\,\mathcal{M}_{\alpha,p}$ (hence $s(A)\geq s(B)$) the infimum in (5.2) is attained by a reverse test (called the “minimal reverse test”) $(\Gamma,a,b)$ with $\mathrm{supp}\,a\supset\mathrm{supp}\,b$. Hence the second inequality in (5.8) holds on $\mathrm{dom}\,\mathcal{M}_{\alpha,p}$.
@@ 5.3 Joint concavity/convexity of trace functions
It is worthwhile to compare condition of Theorem 5.9(2) with the $p=1$ case of (iii) of Theorem 3.2(2). Notice that the former condition is strictly stronger than the latter though they are quite similar.
@@
Note that the ‘only if’ parts of (1) and (2) of Theorem 5.9 are verified by Theorem 5.8. Indeed, if $\mathrm{Tr}\,R_{\alpha,p}$ is jointly concave, then we have $0<\alpha<1$ and $1/p\geq\alpha$ by Theorem 5.8 and Proposition 4.2(1), so we have $1/p\geq 1-\alpha$ too since $R_{\alpha,p}(A,B)=R_{1-\alpha,p}(B,A)$. If $\mathrm{Tr}\,R_{\alpha,p}$ is jointly convex, then we have $\alpha>1$ and $\max\{\alpha/2,\alpha-1\}\leq 1/p\leq\alpha$.
@@
(2) Let $\alpha>1$. Let $A:=\begin{bmatrix}a&0\\ 0&b\end{bmatrix}$ with $a,b>0$, $a\neq b$, and $B:={1\over 2}\begin{bmatrix}1&1\\ 1&1\end{bmatrix}$, and let $\mathcal{E}_{A}$ be the pinching with respect to $A$. It was shown in [47, Lemma 3.17] that $\mathrm{Tr}\,LE_{\alpha}(A,\mathcal{E}_{A}(B))>\mathrm{Tr}\,LE_{\alpha}(A,B)$ for some choice of $a,b$. (Note that $\mathrm{Tr}\,LE_{\alpha}(A,B)$ is denoted as $Q_{\alpha}^{\flat}(B\|A)$ in [47].) Let $U:=\begin{bmatrix}1&0\\ 0&-1\end{bmatrix}$. Since $A={1\over 2}(A+UAU^{*})$, $\mathcal{E}_{A}(B)={1\over 2}(B+UBU^{*})$ and $\mathrm{Tr}\,LE_{\alpha}(A,B)=\mathrm{Tr}\,LE_{\alpha}(UAU^{*},UBU^{*})$, we have
@@
In [47, Theorem 3.6] Mosonyi and Ogawa showed the variational expression
@@
    If $1<\alpha\leq 2$ and $0<p\leq 1$, then for any fixed $B\in\mathbb{M}_{n}^{+}$, $A\mapsto\mathrm{Tr}\,G_{\alpha,p}(A,B)$ is convex on $\{A\in\mathbb{M}_{n}^{+}:s(A)\geq s(B)\}$, $n\in\mathbb{N}$.
@@
(3) The ‘if’ part is a general result for the operator perspective for operator convex functions on $(0,\infty)$ due to [18, 19]. For the ‘only if’ part, a more general and stronger result was shown in [35, Proposition A.1]; see Remark 5.17 below.
@@
Since $1<1-\alpha<0$, note that $x^{1-\alpha}$ ($x>0$) is operator monotone decreasing and operator convex. Hence one has
@@
In view of the proof of [35, Proposition A.1] together with [27, Proposition 3.1], Proposition 5.16(3) is in fact improved in such a way that either if $A\mapsto\mathrm{Tr}\,G_{\alpha,1}(A,B)$ on $\mathbb{M}_{2}^{++}$ for any fixed $B\in\mathbb{M}_{2}^{++}$ or if $B\mapsto\mathrm{Tr}\,G_{\alpha,1}(A,B)$ on $\mathbb{M}_{2}^{++}$ for any fixed $A\in\mathbb{M}_{2}^{++}$ is convex, then $1<\alpha\leq 2$.
@@
for all $A,B\geq 0$ with $s(A)\geq s(B)$. The first inequality gives no restriction (see Proposition 3.13(1)) but the second does the restriction $\alpha p\leq\min\{\alpha,1-\alpha\}$ so that $p\leq\min\bigl\{1,{1-\alpha\over\alpha}\bigr\}$. One can replace $\alpha$ with $1-\alpha$ to have $p\leq\min\bigl\{1,{\alpha\over 1-\alpha}\}$. Therefore, $p\leq\min\bigl\{{1-\alpha\over\alpha},{\alpha\over 1-\alpha}\bigr\}$ must hold. Next, if $\mathrm{Tr}\,\widetilde{S}G_{\alpha,p}$ is jointly concave, then $\mathrm{Tr}\,\widetilde{S}G_{\alpha,p}(A,B)\leq\mathrm{Tr}\,R_{\alpha,1/\alpha}(A,B)$ for all $A,B\geq 0$ with $s(A)\geq s(B)$, which gives $p\leq 1$ by Theorem 3.6(1).
(2) For $\alpha>1$, assume that $\mathrm{Tr}\,SG_{\alpha,p}$ is jointly convex; then by Theorem 5.8(ii), $\mathrm{Tr}\,R_{\alpha,1/\alpha}\leq\mathrm{Tr}\,SG_{\alpha,p}$ on $\mathbb{M}_{2}^{++}\times\mathbb{M}_{2}^{++}$ must hold. But this contradicts Theorem 3.4(2). Assume that $\mathrm{Tr}\,\widetilde{S}G_{\alpha,p}$ is jointly convex. Then we must have $\mathrm{Tr}\,R_{\alpha,1/\alpha}\leq\mathrm{Tr}\,\widetilde{S}G_{\alpha,p}$ on $\mathbb{M}_{2}^{++}\times\mathbb{M}_{2}^{++}$, which contradicts Theorem 3.7(1). ∎
@@ 6 Concluding remarks
(1) In the present paper we have followed the previous paper [33] to discuss the $\alpha$-weighted quasi-geometric type matrix means $R_{\alpha,p}$, $G_{\alpha,p}$, $SG_{\alpha,p}$, $\widetilde{S}G_{\alpha,p}$ and $LE_{\alpha}$ not only for $0<\alpha<1$ but also $\alpha>1$ as defined in Section 2. We are concerned with log-majorizations $\mathcal{M}_{\alpha,p}\prec_{\log}\mathcal{N}_{\alpha,q}$ for pairs $(\mathcal{M},\mathcal{N})$ from $R,G,SG,\widetilde{S}G,LE$ and for $\alpha\in(0,\infty)\setminus\{1\}$, $p>0$. Already known cases are summarized in Theorem 3.1. In Sections 3.13.4 we have examined unknown cases with the aim of finding the necessary and sufficient condition on $p,q,\alpha$ under which $\mathcal{M}_{\alpha,p}(A,B)\prec_{\log}\mathcal{N}_{\alpha,q}(A,B)$ holds for all $A,B>0$. But in many cases we have only a certain necessary condition and/or a certain sufficient condition for that, and the problem is still left open as explained in Problems 3.5, 3.8, 3.12 and 3.20. Especially interesting for us is to characterize $SG_{\alpha,p}\prec_{\log}SG_{\alpha,q}$ and $\widetilde{S}G_{\alpha,p}\prec_{\log}\widetilde{S}G_{\alpha,q}$ (see Problem 3.12).
@@
(2) Another topic we have addressed is the joint concavity/convexity of the trace functions $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ for each $\mathcal{M}$ from $R,G,SG,\widetilde{S}G,LE$. The cases of $\mathrm{Tr}\,R_{\alpha,p}$ ($\alpha>0$) and $\mathrm{Tr}\,G_{\alpha,p}$ ($0<\alpha<1$) are already known [53, 30]. In Section 5 we have aimed at characterizing the joint concavity/convexity of $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ for unknown cases. Our method is based on the close connection between the joint concavity/convexity of geometric type trace functions and the monotonicity under CPTP maps of the generated Rényi type divergences described in Theorem 5.3. Appealing to theory of quantum $\alpha$-Rényi divergences we have obtained a necessary condition for $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ to be jointly concave/convex. The condition is the sandwiched inequalities of $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ with $\mathrm{Tr}\,R_{\alpha,1/\alpha}$ and $\mathrm{Tr}\,G_{\alpha,1}$ (see Theorem 5.8), which are connected with the log-majorizations discussed in Section 3. However, we cannot obtain a sufficient condition for the joint concavity/convexity of $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ in this way, while the necessary condition obtained can be best possible, for example, as exemplified in Remark 5.10. Thus the problem is still left open for $\mathrm{Tr}\,G_{\alpha,p}$ ($\alpha>1$), $\mathrm{Tr}\,SG_{\alpha,p}$ and $\mathrm{Tr}\,\widetilde{S}G_{\alpha,p}$ as mentioned in Problems 5.18 and 5.20.
@@
(3) We have used the Lie–Trotter–Kato product formula (Theorem 2.2 and [33, Theorem 2.3]) in the proofs of (1), (4) and (5) of Corollary 3.18 and Theorem 5.12(1) for example. In fact, in view of Proposition 2.1(2) we may assume that $A,B>0$ in the proofs of those, in which case the proof of Theorem 2.2 becomes much simpler. Note however that in the proof of Proposition 2.1(2) we have used [33, Lemma A.3] that is essential in the proof of [33, Theorem 2.3].
@@
(5) It is well known [18, 19, 35, 37] that the operator perspective

### v3

@@ 5.1 Joint concavity/convexity and monotonicity
The next theorem was shown in [32, Theorem 3.2] in a more general form by a complex function method using Pick functions (called Epstein’s method).
@@
In particular, $\mathrm{Tr}\,\mathcal{A}_{\alpha,p}$, $\mathrm{Tr}\,\mathcal{H}_{\alpha,p}$ and $\mathrm{Tr}\,G_{\alpha,p}$ are jointly concave for $0<\alpha\leq 1$ and $0<p\leq 1$. More specified results on the joint concavity/convexity of $\mathrm{Tr}\,\mathcal{A}_{\alpha,p}$ and $\mathrm{Tr}\,\mathcal{H}_{\alpha,p}$ are known as follows (see [8, 15] and [32, p. 1583]):
@@
Let $0<\alpha<1$, $p>0$, and let $n\in\mathbb{N}$ with $n\geq 2$ be arbitrarily fixed. Then the following hold:
@@ 5.2 Quantum Rényi type divergences
where a POVM on $\mathbb{C}^{n}$ is a family $(M_{i})_{i=1}^{k}\subset\mathbb{M}_{n}^{+}$ with $\sum_{i=1}^{k}M_{i}=I_{n}$ (i.e., a $k$-outcome positive operator-valued measure). Furthermore, the regularized measured $\alpha$-Rényi divergence is defined by
@@
we have the second inequality in (5.5). If, in addition, $D_{\alpha}^{\mathrm{q}}$ is additive under tensor products, then by (5.5) with definition (5.4) we furthermore have
@@
Let $p>0$. Either when $\mathcal{M}\in\{R,G,LE\}$ and $\alpha\in(0,\infty)\setminus\{1\}$, or when $\mathcal{M}\in\{SG,\widetilde{S}G\}$ and $\alpha>1$, the function $D^{\mathcal{M}_{\alpha,p}}$ defined in (5.7) is a quantum $\alpha$-Rényi type divergence in the sense of Definition 5.4, and is additive under tensor products.
@@
    ([48, Remark II.6]) When $\alpha>2$, $D_{\alpha}^{\max}(B\|A)\leq D^{G_{\alpha,1}}(B\|A)$ and $D_{\alpha}^{\max}\neq D^{G_{\alpha,1}}$.
@@
    ([49, Theorem 3.7], [29]) When $1/2\leq\alpha<\infty$,
@@
For any $\mathcal{M}_{\alpha,p}$ in Lemma 5.5, since $\mathcal{M}_{\alpha,p}(a,b)=a^{1-\alpha}b^{\alpha}$ in $a,b\in(0,\infty)$ is not convex when $0<\alpha<1$ and not concave when $\alpha>1$, we may consider only the joint concavity of $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ when $0<\alpha<1$, and the joint convexity of $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ when $\alpha>1$. Below we will simply write $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ is jointly concave/convex” if it is so on the domain of $\mathcal{M}_{\alpha,p}$, equivalently, if it is so on $\mathbb{M}_{n}^{++}\times\mathbb{M}_{n}^{++}$ for any $n\in\mathbb{N}$ (thanks to Proposition 2.1(2)).
@@
for all $(A,B)\in\mathrm{dom}\,\mathcal{M}_{\alpha,p}$, $B\neq 0$. When $\mathcal{M}\in\{SG,\widetilde{S}G\}$ and $0<\alpha<1$, even though $D^{\mathcal{M}_{\alpha,p}}$ is not precisely a quantum $\alpha$-Rényi type divergence in Definition 5.4, we can still have (5.8) for all $(A,B)\in\mathrm{dom}\,\mathcal{M}_{\alpha,p}$. In fact, Theorem 5.3 implies that $D^{\mathcal{M}_{\alpha,p}}$ is monotone under CPTP maps with restriction to $\mathrm{dom}\,\mathcal{M}_{\alpha,p}$. Hence by Lemma 5.6 the first inequality in (5.8) holds on $\mathrm{dom}\,\mathcal{M}_{\alpha,p}$. Also, note (see [47] and the proof of [37, Proposition 4.1]) that for any $(A,B)\in\mathrm{dom}\,\mathcal{M}_{\alpha,p}$ (hence $s(A)\geq s(B)$) the infimum in (5.2) is attained by a reverse test (called the “minimal reverse test”) $(\Gamma,a,b)$ with $\mathrm{supp}\,a\supset\mathrm{supp}\,b$. Hence the second inequality in (5.8) holds on $\mathrm{dom}\,\mathcal{M}_{\alpha,p}$.
@@ 5.3 Joint concavity/convexity of trace functions
It is worthwhile to compare the condition of Theorem 5.9(2) with the $p=1$ (and $q$ replaced with $p$) case of (iii) of Theorem 3.2(2). Notice that the former condition is strictly stronger than the latter though they are quite similar.
@@
Note that the ‘only if’ parts of (1) and (2) of Theorem 5.9 are verified by Theorem 5.8. Indeed, if $\mathrm{Tr}\,R_{\alpha,p}$ is jointly concave, then we have $0<\alpha<1$ and $1/p\geq\alpha$ by Theorem 5.8 and Proposition 4.2(1), so we have $1/p\geq 1-\alpha$ too since $R_{\alpha,p}(A,B)=R_{1-\alpha,p}(B,A)$. If $\mathrm{Tr}\,R_{\alpha,p}$ is jointly convex, then we have $\alpha>1$ and $\max\{\alpha/2,\alpha-1\}\leq 1/p\leq\alpha$ by Proposition 4.2(1) and Theorem 3.2(2) (with Remark 2.4(4)).
@@
(2) Let $\alpha>1$. Let $A:=\begin{bmatrix}a&0\\ 0&b\end{bmatrix}$ with $a,b>0$, $a\neq b$, and $B:={1\over 2}\begin{bmatrix}1&1\\ 1&1\end{bmatrix}$, and let $\mathcal{E}_{A}$ be the pinching with respect to $A$. It was shown in [50, Lemma 3.17] that $\mathrm{Tr}\,LE_{\alpha}(A,\mathcal{E}_{A}(B))>\mathrm{Tr}\,LE_{\alpha}(A,B)$ for some choice of $a,b$. (Note that $\mathrm{Tr}\,LE_{\alpha}(A,B)$ is denoted as $Q_{\alpha}^{\flat}(B\|A)$ in [50].) Let $U:=\begin{bmatrix}1&0\\ 0&-1\end{bmatrix}$. Since $A={1\over 2}(A+UAU^{*})$, $\mathcal{E}_{A}(B)={1\over 2}(B+UBU^{*})$ and $\mathrm{Tr}\,LE_{\alpha}(A,B)=\mathrm{Tr}\,LE_{\alpha}(UAU^{*},UBU^{*})$, we have
@@
In [50, Theorem 3.6] Mosonyi and Ogawa showed the variational expression
@@
    If $1<\alpha\leq 2$ and $0<p\leq 1$, then for any fixed $B\in\mathbb{M}_{n}^{+}$ with any $n\in\mathbb{N}$, $A\mapsto\mathrm{Tr}\,G_{\alpha,p}(A,B)$ is convex on $\{A\in\mathbb{M}_{n}^{+}:s(A)\geq s(B)\}$.
@@
(3) The ‘if’ part is a general result for the operator perspective for operator convex functions on $(0,\infty)$ due to [18, 19]. For the ‘only if’ part, a more general and stronger result was shown in [37, Proposition A.1]; see Remark 5.17 below.
@@
Since $1\leq 1-\alpha<0$, note that $x^{1-\alpha}$ ($x>0$) is operator monotone decreasing and operator convex. Hence one has
@@
In view of the proof of [37, Proposition A.1] together with [27, Proposition 3.1], Proposition 5.16(3) is in fact improved in such a way that either if $A\mapsto\mathrm{Tr}\,G_{\alpha,1}(A,B)$ on $\mathbb{M}_{2}^{++}$ for any fixed $B\in\mathbb{M}_{2}^{++}$ is convex or if $B\mapsto\mathrm{Tr}\,G_{\alpha,1}(A,B)$ on $\mathbb{M}_{2}^{++}$ for any fixed $A\in\mathbb{M}_{2}^{++}$ is convex, then $1<\alpha\leq 2$.
@@
for all $A,B\geq 0$ with $s(A)\geq s(B)$. The first inequality gives no restriction (see Proposition 3.13(1)), but by Theorem 3.3(1) the second does the restriction $\alpha p\leq\min\{\alpha,1-\alpha\}$ so that $p\leq\min\bigl\{1,{1-\alpha\over\alpha}\bigr\}$. One can replace $\alpha$ with $1-\alpha$ to have $p\leq\min\bigl\{1,{\alpha\over 1-\alpha}\}$. Therefore, $p\leq\min\bigl\{{1-\alpha\over\alpha},{\alpha\over 1-\alpha}\bigr\}$ must hold. Next, if $\mathrm{Tr}\,\widetilde{S}G_{\alpha,p}$ is jointly concave, then $\mathrm{Tr}\,\widetilde{S}G_{\alpha,p}(A,B)\leq\mathrm{Tr}\,R_{\alpha,1/\alpha}(A,B)$ for all $A,B\geq 0$ with $s(A)\geq s(B)$, which gives $p\leq 1$ by Theorem 3.6(1).
(2) For $\alpha>1$, assume that $\mathrm{Tr}\,SG_{\alpha,p}$ is jointly convex; then by Theorem 5.8(ii), $\mathrm{Tr}\,R_{\alpha,1/\alpha}\leq\mathrm{Tr}\,SG_{\alpha,p}$ on $\mathbb{M}_{2}^{++}\times\mathbb{M}_{2}^{++}$ must hold. But this contradicts Theorem 3.4(2). Assume that $\mathrm{Tr}\,\widetilde{S}G_{\alpha,p}$ is jointly convex. Then we must have $\mathrm{Tr}\,\widetilde{S}G_{\alpha,p}\leq\mathrm{Tr}\,G_{\alpha,1}$ on $\mathbb{M}_{2}^{++}\times\mathbb{M}_{2}^{++}$, which contradicts Proposition 3.14(1). ∎
@@ 6 Concluding remarks
(1) In the present paper we have followed the previous paper [35] to discuss the $\alpha$-weighted quasi-geometric type matrix means $R_{\alpha,p}$, $G_{\alpha,p}$, $SG_{\alpha,p}$, $\widetilde{S}G_{\alpha,p}$ and $LE_{\alpha}$ not only for $0<\alpha<1$ but also $\alpha>1$ as defined in Section 2. We are concerned with log-majorizations $\mathcal{M}_{\alpha,p}\prec_{\log}\mathcal{N}_{\alpha,q}$ for pairs $(\mathcal{M},\mathcal{N})$ from $\{R,G,SG,\widetilde{S}G,LE\}$ and for $\alpha\in(0,\infty)\setminus\{1\}$, $p>0$. Already known cases are summarized in Theorem 3.1. In Sections 3.13.4 we have examined unknown cases with the aim of finding the necessary and sufficient condition on $p,q,\alpha$ under which $\mathcal{M}_{\alpha,p}(A,B)\prec_{\log}\mathcal{N}_{\alpha,q}(A,B)$ holds for all $A,B>0$. But in many cases we have only a certain necessary condition and/or a certain sufficient condition for that, and the problem is still left open as explained in Problems 3.5, 3.8, 3.12 and 3.20.
(2) A natural and interesting question is to find $\mathcal{M},\mathcal{N}$ from $\{R,G,SG,\widetilde{S}G,LE\}$ and $\alpha,p,q$ such that the following (a) fails to hold but (b) holds: (a) $\mathcal{M}_{\alpha,p}(A,B)\prec_{\log}\mathcal{N}_{\alpha,q}(A,B)$ for all $A,B\in\mathbb{M}_{n}^{++}$; (b) $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}(A,B)\leq\mathrm{Tr}\,\mathcal{N}_{\alpha,q}(A,B)$ for all $A,B\in\mathbb{M}_{n}^{++}$. Note that (a) and (b) are equivalent when $n=2$ by Remark 2.4(4), and are also equivalent for all $n$ in certain cases of $(\mathcal{M}_{\alpha,p},\mathcal{N}_{\alpha,q})$ due to Theorems 3.2, 3.3, 3.6 and Proposition 4.2.
(3) In a recent paper [55] Seo, Wada and Yamazaki have, appealing to the grand Furuta inequality [22], shown Proposition 3.10(1) and improved Proposition 3.10(2) in such a way that if $0<\alpha\leq 1/2$ and $p/q\leq 2\alpha$, then $\widetilde{S}G_{\alpha,p}(A,B)\prec_{\log}\widetilde{S}G_{\alpha,q}(A,B)$ for all $A,B>0$. Moreover, it has been shown in [55, Theorem 3.13] that for some $\alpha\in(1/2,1)$ and for any $p\in(0,1)$ there exist $A,B\in\mathbb{M}_{2}^{++}$ such that $\widetilde{S}G_{\alpha,p}(A,B)\not\prec_{\log}\widetilde{S}G_{\alpha,1}(A,B)$. Indeed, in view of Theorem 3.19 (with Theorem 2.2), the same holds true for any $\alpha\in(1/2,1)$. From this and Proposition 3.11(3) we conclude that $\widetilde{S}G_{\alpha,p}\prec_{\log}\widetilde{S}G_{\alpha,q}$ fails to hold for both $p<q$ and $p>q$ when $1/2<\alpha<1$. Thus, concerning Problem 3.12 on $\widetilde{G}_{\alpha,p}$ for $0<\alpha<1$, the remaining is whether or not $\widetilde{S}G_{\alpha,p}\prec_{\log}\widetilde{S}G_{\alpha,q}$ holds if $0<\alpha<1/2$ and $0<p<q$, extending the above mentioned case $p/q\leq 2\alpha$.
@@
(4) Another topic we have addressed is the joint concavity/convexity of the trace functions $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ for each $\mathcal{M}$ from $\{R,G,SG,\widetilde{S}G,LE\}$. The cases of $\mathrm{Tr}\,R_{\alpha,p}$ ($\alpha>0$) and $\mathrm{Tr}\,G_{\alpha,p}$ ($0<\alpha<1$) are already known [57, 32]. In Section 5 we have aimed at characterizing the joint concavity/convexity of $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ for unknown cases. Our method is based on the close connection between the joint concavity/convexity of geometric type trace functions and the monotonicity under CPTP maps of the generated Rényi type divergences described in Theorem 5.3. Appealing to theory of quantum $\alpha$-Rényi divergences we have obtained a necessary condition for $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ to be jointly concave/convex. The condition is the sandwiched inequalities of $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ with $\mathrm{Tr}\,R_{\alpha,1/\alpha}$ and $\mathrm{Tr}\,G_{\alpha,1}$ (see Theorem 5.8), which are connected with the log-majorizations discussed in Section 3. However, we cannot obtain a sufficient condition for the joint concavity/convexity of $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ in this way, while the necessary condition obtained can be best possible, for example, as exemplified in Remark 5.10. Thus the problem is still left open for $\mathrm{Tr}\,G_{\alpha,p}$ ($\alpha>1$), $\mathrm{Tr}\,SG_{\alpha,p}$ and $\mathrm{Tr}\,\widetilde{S}G_{\alpha,p}$ as mentioned in Problems 5.18 and 5.20.
@@
(5) We have used the Lie–Trotter–Kato product formula (Theorem 2.2 and [35, Theorem 2.3]) in the proofs of (1), (4) and (5) of Corollary 3.18 and Theorem 5.12(1) for example. In fact, in view of Proposition 2.1(2) we may assume that $A,B>0$ in the proofs of those, in which case the proof of Theorem 2.2 becomes much simpler. Note however that in the proof of Proposition 2.1(2) we have used [35, Lemma A.3] that is essential in the proof of [35, Theorems A.1 and 2.3].
@@
(7) It is well known [43, 19, 18, 37, 39] that the operator perspective